The Pristine River has two polluting firms on its banks. European Industrial and Creative Chemicals each dump 100 tonnes of effluent into the river each year. The cost of reducing effluent emissions per tonne equals €300 for European Industrial and €200 for Creative. The government wants to reduce overall pollution from 200 tonnes to 40 tonnes per year. a. If the government knew the cost of reduction for each firm, what reductions would it impose to reach its overall goal? What would be the cost to each firm and the total cost to the firms together? b. In a more typical situation, the government would not know the cost of pollution reduction at each firm. If the government decided to reach its overall goal by imposing uniform reductions on the firms, calculate the reduction made by each firm, the cost to each firm and the total cost to the firms together.


1
Expert's answer
2021-12-20T10:20:36-0500

a. If the government knew the cost of reduction for each firm, then it would impose the larger reduction on the firm with lower cost of reducing (Creative Chemicals) to reach its overall goal.

If Creative Chemicals decreases production by 100 tones at cost of 200×100 = €20,000 and European Industrial decreases production by 60 tones at cost 300×60 = €18,000, then the total cost to the firms together is €38,000.

b. If the government decided to reach its overall goal by imposing uniform reductions on the firms, then each firm would reduce production by 80 tones, the cost to Creative Chemicals would be 80×200 = €16,000 and the cost to European Industrial 80×300 = €24,000, the total cost to the firms together would be €40,000.